Report on NIRSPEC observing nights of 14-16 December 2000

Summary Report on the three nights of Gemini/ NIRSPEC 2000B time

The three Gemini/NIRSPEC observing nights on Keck II in Semester 2000B were December 14/15, 15/16, and 16/17. Clouds covered most of the sky through the first night, but generally decreased through the night, with large clear patches present by dawn. Due to the clouds half of the night was completely lost and the other half was compromised to the extent that only two programs could be observed, with considerably longer integrations than had been requested. On the second and third nights skies were clear and, apart from about one hour lost to a technical problem (transputer hangup) and a few other minor problems, observing went very smoothly. Seeing at K was ~0.8 arcsec on the first night, ~0.6 arcsec on the second night and ~0.4 arcsec on the third night.

While atttempting to carry out programs in the thermal IR (3-5um) we noticed a problem - the appearance of numerous transient ambient temperature point sources which were not astronomical. Many of these were of comparable brightness to the science targets. This problem was called to the attention of Keck staff and identified as due to splotches and marks on the NIRSPEC window. Although we continued observing at thermal IR wavelengths during the run, it is clear that the quality of the thermal IR spectra are severely compromised. PIs of these programs have been notified of the problem.

Apart from the above, NIRSPEC and the telescope functioned more accurately and efficiently than in the previous semester. The instrument has a new array with many fewer hot pixels and only an occasional blotch. The new array has the same persistence problem as its predecessor. Although we believe that in most cases we were able to avoid or work around the effect, some data may be compromised. Because the instrument has no shutter, observations of (bright) calibration stars almost inevitably lead to overexposing the array during set-up. It takes about 30-60 minutes for the effect to fade away.

The following are programs that were attempted, along with brief statements of the results. Within the next two weeks we will be making the data available to the PI's of the proposals via ftp.

075csk (lensed ERO - Kochanek) - K band spectrum obtained on second night, believed to be good quality.

079k_n (Methane in L Dwarfs) - L band spectra obtained of both targets on second night; compromised by dust and blemishes on dewar window.

082s_g (Dust absorption in Lyman Break galaxies - Gwyn) - one target observed in the K band through clouds at the end of the first night and in the H band on the second night.

084cgt (Rotation of Brown Dwarfs - Tinney) - three targets observed on third night; spectra appear to be good quality.

085s_b (O abundance - Balachandran) - one target observed on the third night, compromised by dust and blemishes on dewar window.

087j_m (O abundances - Melendez) - Two targets observed at the start of the first night.

096txt (I Zw 18 - Thuan) - long and short J spectra obtained on third night; possibly compromised by persistence.

098s_v (Starburst at z=0.4 - Veilleux) J and H band spectra obtained on night 2; continuum easily seen.

099jaf (Globular Clusters in M31 - Frogel) - K band spectra obtained of both targets on night 3; believed to be good quality.

101j_c (Gaps in Disks - Carr) - both targets observed on night 3; integration time reduced for fainter target. Data compromised by dust and blemishes on dewar window.

104bjh (Proto-planetary Nebulae - Hrivnak) - three targets measured on night 2; spectrum of the third curtailed by transputer problem. Data compromised by blemishes on dewar window, but targets are bright so data may be useful

106g_r (Clusters in M82 - Rieke) - spectrum obtained on night 3. One hour integration only, but s/n looks good.

110m_t (SFR at z>2 - Takamiya) - One target observed on third night.

[Science Operations home] [NIRSPEC home]

Last update March 22, 2001; Tom Geballe